OpenStudy (anonymous):

Someone Please Help me step by step! Solve the system by graphing (no need to submit your graph here) then write the solution as an ordered pair). 7. y = -1/2x - 1 y = 1/2x - 3 Solution: ( , )

OpenStudy (anonymous):

find x-intercept and y-intercept of both the lines sketch them their joining point is ur answer

OpenStudy (lukebluefive):

The equations are in slope-intercept form: \[y = mx + b\] Here, m is the slope: \[slope = \frac{rise}{run}\] Now, to draw the lines, first find where they intercept each axis. This isn't so bad, since you just set y=0 to get the x-intercept: \[0 = mx + b\] Subtract b from both sides: \[-b = mx\] Then, solve for x: \[-\frac{b}{m} = x\] Then, set x=0 to get the y-intercept: \[y = m \times 0 + b\]

OpenStudy (lukebluefive):

After that, you should be able to draw both lines and find where they intersect.

OpenStudy (lukebluefive):

The y-intercepts will be points of the form: \[(x, 0)\] While the x-intercepts will be of the form: \[(0, y)\]

OpenStudy (lukebluefive):

Where x and y are the values you got from setting the other variable to zero, as I explained above.

OpenStudy (anonymous):

Thank you so much, You explained that very well! @LukeBlueFive

OpenStudy (lukebluefive):

@JessNicole You're very welcome!